this is my problem. 19+x/2=4 because it's division i dont know how to solve. i dont understand to move the 19 over and move the 4 on the other side of the = sign. please help.
signed
below grade level.


Found 2 solutions by checkley77, Masaries7:
Answer by checkley77(7072)

19+x/2=4 subtract 19 from both sides.
x/2=4-19
x/2=-15 cross multiply.
x=2*-15
x=-30 ans.
Proof:
19-30/2=4
19-15=4
4=4

Answer by Masaries7(3)

Here's the problem restated:
Step 1. Isolate the variable by it's self

on the left hand side the two "19"'s cancel becoming...

"4-19" equals "-15" so rewriting

Step 2. Reverse whatever process is being done to "x" if "x" is being multiplied, you divide, if it's being divided, you multiply. etc

multiply both sides by two

The left hand side with "x" is being divided by "2" and multiplied by "2" these effects cancel each other out.

multiply the right hand side

To check our answer we input the solution we concluded which was "-30" into the original problem which was




Since we arrived at a mathematical truth (that "4" is indeed "4") the solution we concluded is indeed correct. It's my wish that you will be able to understand any type of problem of this manner. Please let me know if you have any further questions regarding this problem or any other math problem that's a puzzler for you.